OpenStudy (anonymous):

The numbers of seats in the first 12 rows of a high-school auditorium form an arithmetic sequence. The first row has 9 seats. The second row has 11 seats. a) Write a recursive formula to represent the sequence. b) Write an explicit formula to represent the sequence.

Directrix (directrix):

The following is not my work although I checked it. a) Recursive formulas describe the relationship between a term and the previous one in the same sequence. In this case, a1=9 and a2=11 (1 and 2 are the rows). So an= a(n-1)+2 (difference between 9 and 11 is 2)

Directrix (directrix):

b) Explicit formulas allow you to calculate an irrespective of whether you have the previous or latter term or not. In this case: an= 2n+7 (if you plug 1, you get 2(1)+7 = 9; all you need is the row number and you can get the number of seats using the explicit formula. c) a(12)= 2(12) + 7 = 31 seats Solution courtesy of unknown solver at this location: http://tinyurl.com/bhcg2gt